# Congestion-Approximators from the Bottom Up

### v2

@@ 5.2 Fair Cut/Flow Algorithm
Given a congestion-approximator, the most convenient min-cut/max-flow algorithm is the fair cut/flow algorithm of [25]. We state the definition of a fair cut/flow and then cite the main result of [25].
@@
Theorem 5.9 (Fair cut/flow algorithm [25]).
@@ 5.3 Cut-Matching Game and Trimming
We follow the cut-matching game treatment in [34]: either find a “balanced” cut of small capacity, or ensure that a “large” part of the graph mixes with low congestion. The following lemma is similar to Theorem 2.2 of [34] with one key difference: there is no built-in flow subroutine, so the algorithm makes black-box calls to the fair cut/flow algorithm of Theorem 5.10.
In past work [33, 1], the analysis of the cut-matching game for capacitated graphs has only been sketched, referencing the fact that a capacitated graph can be modelled by an uncapacitated graph with parallel edges (at a cost). For completeness, we provide a full proof of this capacitated case in Appendix A.
@@
Property (3) asserts that either an approximately balanced cut is found, or the vertex weighting $\mathbf{d}|_{A\setminus R}$ mixes with low congestion (with high probability). In our algorithm for Theorem 5.1, we actually want the weighting $\mathbf{d}_{A\setminus R}+\deg_{\partial R}$ to mix in the second case. To guarantee this stronger property, we augment the set $R$ into $R\cup B$ through one additional call to the fair cut/flow algorithm of Theorem 5.10. The algorithm is similar to the flow-based expander trimming procedure in [34]. For completeness, we defer the algorithm and proof to Appendix B. Note that the setting, including assumption 1, is the same as Theorem 5.11.
@@ 5.4 Clustering Algorithm
For any recursive call $A^{\prime}\subseteq A$, we have $\mathbf{d}^{\prime}(A^{\prime})\leq(1-\frac{1}{24T})\mathbf{d}(A)$ where $\mathbf{d}^{\prime}=\textup{deg}_{\partial\mathcal{P}_{L}\cup\partial A^{\prime}}|_{A^{\prime}}$.
@@
$\leq\frac{7}{6}\cdot\frac{1}{2}\mathbf{d}(A)+\frac{1}{36}\mathbf{d}(A)+\frac{1}{6}\mathbf{d}(A)$
$=\frac{7}{9}\mathbf{d}(A)$
$\leq\left(1-\frac{1}{24T}\right)\mathbf{d}(A),$
@@
For a given recursive call $A$, define its recursion depth inductively as follows: the initial call $A\leftarrow V$ has depth 00, and given a recursive call $A$ of depth $d$, all of its recursive calls have depth $d+1$. By Claim 5.14, the value of $\mathbf{d}(A)$ decreases multiplicatively by factor $1/(24T)$ on each recursive call, so the maximum recursion depth is $O(T\log(nW))$.
@@ 6 Approximate Maximum Flow
From Theorem 5.1 and Claim 5.2, we obtain an algorithm that constructs a congestion-approximator of quality $O(\log^{10}(nW))$ in $\tilde{O}(m)$ time. Recall that Sherman’s framework [36, 37] translates a congestion-approximator of quality $\alpha$ to a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}\alpha m)$. Thus, for any parameter $\epsilon>0$, we obtain a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}m)$.
@@ Appendix A Cut-Matching Game
Our setup resembles Appendix B of [34] with a few minor changes. In particular, since our flow routine is more restrictive, we have to adapt the algorithm to handle our flow outputs.
We begin with notation from [34]. For simplicity, we avoid working with the subdivision graph in [33, 34]. Define a $A$-commodity flow as a multi-commodity flow where each vertex $v\in A$ is the source of quantity $\mathbf{d}(v)$ of its distinct flow commodity. Only for analysis, we consider a $A\times A$ flow-matrix $\mathbf{F}\in\mathbb{R}^{A\times A}_{\geq 0}$ which encodes information about a $A$-commodity flow. We say that $\mathbf{F}$ is routable with congestion $c$ if there exists a $A$-commodity flow $f$ such that, simultaneously for all $u,v\in A$, we have that $u$ can send quantity $\mathbf{F}(u,v)$ of its own commodity to $v$, and the amount of flow through each edge is at most $c$.
The algorithm initializes flow-matrix $\mathbf{F}_{0}\in\mathbb{R}^{A\times A}_{\geq 0}$ as the diagonal matrix with value $\mathbf{d}(v)$ on entry $\mathbf{F}(v,v)$. Trivially, $\mathbf{F}$ is routable with zero congestion. The algorithm initializes $A_{0}=A$ and $R_{0}=\emptyset$, and then proceeds for at most $T=O(\log^{2}(nW))$ rounds. For each round $t\in[T]$, the algorithm implicitly updates $\mathbf{F}_{t-1}$ to $\mathbf{F}_{t}$ such that it is routable with congestion at most $t/\phi$. The operation for implicitly updating $\mathbf{F}_{t-1}$ will be described explicitly later on, but we ensure that row sums do not change from $\mathbf{F}_{t-1}$ to $\mathbf{F}_{t}$, i.e., there is always $\mathbf{d}(v)$ total quantity of each commodity $v\in A$ spread among the vertices. For each round $t\in[T]$, the algorithm (explicitly) finds a partition of $A_{t-1}$ into $A_{t}^{\ell},A_{t}^{r}$, and then computes
@@
    A (possibly empty) set $S_{t}\subseteq A$ satisfying $\delta_{G[A]}S_{t}\leq\phi\mathbf{d}(S\cap A_{t-1})+\frac{\phi}{6T^{2}}\mathbf{d}(A)$ and $\mathbf{d}(S_{t})\leq\mathbf{d}(A)/3$, and
@@
    A (possibly empty) flow $f_{t}$ from $A_{t}^{\ell}\setminus S_{t}$ to $A_{t}^{r}\setminus S_{t}$ such that each vertex $v\in A_{t}^{\ell}\setminus S_{t}$ is the source of at least $\mathbf{d}(v)/12$ flow, and each vertex $v\in A_{t}^{r}\setminus S_{t}$ is the sink of at most $\mathbf{d}(v)$ flow. The flow $f$ has congestion $1/\phi$.
@@
For any round $t$, we have $\delta_{G[A]}R_{t}\leq\phi\mathbf{d}(R_{t})+\frac{\phi}{6T}\mathbf{d}(A)$ and $\mathbf{d}(R_{t})\leq\mathbf{d}(A)/2$.
@@
For the second statement, consider the round $t$ with $\mathbf{d}(R_{t})\geq\mathbf{d}(A)/(6T)$, if it exists, at which point the algorithm terminates. (If there is no such $t$, then we are done.) We have $\mathbf{d}(R_{t-1})<\mathbf{d}(A)/(6T)$ and $\mathbf{d}(R_{t-1}\cup S_{t})=\mathbf{d}(R_{t})\geq\mathbf{d}(A)/(6T)$, and the final set $S_{t}$ satisfies $\mathbf{d}(S_{t})\leq\mathbf{d}(A_{t-1})/3\leq\mathbf{d}(A)/3$. It follows that the new set $R_{t}$ satisfies
$\mathbf{d}(R_{t})=\mathbf{d}(R_{t-1}\cup S_{t-1})\leq\mathbf{d}(R_{t-1})+\mathbf{d}(S_{t})\leq\frac{\mathbf{d}(A)}{6T}+\frac{\mathbf{d}(A)}{3}\leq\frac{\mathbf{d}(A)}{2},$
concluding the second statement and the proof. ∎
@@
$\boldsymbol{\mu}_{t}=\frac{\sum_{u\in A_{t}}\mathbf{F}_{t}(u)}{\mathbf{d}(A_{t})}=\arg\min_{\boldsymbol{\mu}\in\mathbb{R}^{A}}\sum_{u\in A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}\right\rVert_{2}^{2}$ (2)
is the weighted average of the flow-vectors in $A_{t}$, which is also the minimizer of $\psi(t)$ when treated as a function of $\psi(t)$. The latter fact can be verified separately for each coordinate $v\in A$; namely,
@@ A.2 Computing the Set and Flow
In this subsection, we describe a round of the algorithm in detail, following Lemma B.3 of [34] with some minor changes. Recall that on each iteration $t\in[T]$, the algorithm first computes a partition of $A_{t-1}$ into $A_{t}^{\ell},A_{t}^{r}$, and then computes a set $S_{t}$ and flow $f_{t}$ satisfying certain properties. The algorithm then updates $A_{t}\leftarrow A_{t-1}\setminus S_{t}$ and $R_{t}\leftarrow R_{t-1}\cup S_{t}$. The choice of partition $A_{t}^{\ell},A_{t}^{r}$ will depend on an implicitly represented flow-matrix $\mathbf{F}_{t}$ that is useful for the analysis.
@@ A.2.1 Constructing the partition
Lemma A.5.
@@
    $\mathbf{d}(A_{t}^{\ell})\leq\mathbf{d}(A_{t-1})/2$ and $\mathbf{d}(A_{t}^{r})\geq\mathbf{d}(A_{t-1})/2$, and
@@
    $\sum_{v\in A_{t}^{\ell}}\mathbf{d}(v)\cdot(\mathbf{p}(v)-\eta)^{2}\geq\frac{1}{2}\sum_{v\in A_{t-1}}\mathbf{d}(v)\cdot(\mathbf{p}(v)-\bar{\mu}_{t-1})^{2}$.
Proof.
Sort the vertices $v\in A_{t-1}$ by their value of $\mathbf{p}(v)$ in ascending order, and let the sorted list be $v_{1},v_{2},\ldots,v_{|A_{t-1}|}$. Let $i\geq 1$ be the smallest integer such that $\mathbf{d}(\{v_{1},v_{2},\ldots,v_{i}\})\geq\mathbf{d}(A_{t-1})/2$. Define the vertex sets $S_{1}=\{v_{1},v_{2},\ldots,v_{i}\}$ and $S_{2}=\{v_{i},v_{i+1},\ldots,v_{|A_{t-1}|}\}$. By our construction of $i$, we have $\mathbf{d}(S_{j}\setminus\{v_{i}\})<\mathbf{d}(A_{t-1})/2$ for both $j\in\{1,2\}$. Since $S_{1}\cup S_{2}=A_{t-1}$, we have
$\sum_{v\in S_{1}}\mathbf{d}(v)\cdot(\mathbf{p}(v)-\eta)^{2}+\sum_{v\in S_{2}}\mathbf{d}(v)\cdot(\mathbf{p}(u)-\eta)^{2}\geq\sum_{u\in A_{t-1}}\mathbf{d}(v)\cdot(\mathbf{p}(u)-\eta)^{2},$
so pick $j\in\{1,2\}$ such that
$\sum_{v\in S_{j}}\mathbf{d}(v)\cdot(\mathbf{p}(u)-\eta)^{2}\geq\frac{1}{2}\sum_{u\in A_{t-1}}\mathbf{d}(v)\cdot(\mathbf{p}(u)-\eta)^{2}.$
Define $\eta=p_{v_{i}}$ so that
$\sum_{v\in S_{j}\setminus\{i\}}\mathbf{d}(v)\cdot(\mathbf{p}(u)-\eta)^{2}=\sum_{v\in S_{j}}\mathbf{d}(v)\cdot(\mathbf{p}(u)-\eta)^{2}\geq\frac{1}{2}\sum_{u\in A_{t-1}}\mathbf{d}(v)\cdot(\mathbf{p}(u)-\eta)^{2}.$
It follows that we can take $A_{t}^{\ell}=S_{j}\setminus\{i\}$ and $A_{t}^{r}=A_{t-1}\setminus A_{t}^{\ell}$, which satisfies all three properties of the lemma. ∎
We conclude this section with a lemma about the behavior of projections onto a random unit vector. Since the lemma is standard in high-dimensional geometry, we omit the proof and refer to Lemma 3.4 of [18].
Lemma A.6 (Lemma 3.4 of [18]).
@@
with probability at least $1-n^{-c/4}$. In particular, there exists a constant $C>0$ such that for any pair $u,v\in A$,
$\mathbb{E}[(\mathbf{p}(u)-\mathbf{p}(v))^{2}]\leq\frac{C\log n}{n}\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\frac{\mathbf{F}_{t-1}(v)}{\mathbf{d}(v)}\right\rVert_{2}^{2}.$
@@ A.2.2 Max-flow/min-cut call
$A\leftarrow A,\,\epsilon\leftarrow\frac{1}{18T^{2}},\,\gamma\leftarrow\frac{\epsilon\phi}{2},\,\beta\leftarrow\max\{1,(12\phi+\epsilon\gamma)(\kappa+2)\},\,\,\mathbf{s}\leftarrow\phi\mathbf{d}|_{A_{t}^{\ell}}+\epsilon\phi\mathbf{d},\,\mathbf{t}\leftarrow 12\phi\mathbf{d}|_{A_{t}^{r}},$ (3)
@@
From Theorem 5.10, we obtain an $(1+\epsilon)$-approximate fair cut/flow pair $(S,f)$. The algorithm sets $S_{t}\leftarrow S\setminus\{s,x\}$, which satisfies the condition below as required by step i.
@@
$\delta_{G[A]}S_{t}\leq\phi\mathbf{d}(S\cap A_{t-1})+3\epsilon\phi\mathbf{d}(A)$ and $\mathbf{d}(S_{t})\leq\mathbf{d}(A)/3$,
@@
    The edges incident to $s$, which have total capacity $c_{H}(\{s\},V_{H}\setminus S)=\mathbf{s}(A\setminus S_{t})\geq\phi\mathbf{d}(A_{t}^{\ell}\setminus S_{t})$, and
@@
$\delta_{H}S\geq\delta_{G[A]}S_{t}+\phi\mathbf{d}(A_{t}^{\ell}\setminus S_{t}).$
@@
$\delta_{H}S\leq(1+\epsilon)\mathbf{s}(A)=(1+\epsilon)(\phi\mathbf{d}(A_{t}^{\ell})+\epsilon\phi\mathbf{d}(A))\leq(1+\epsilon)\phi\mathbf{d}(A_{t}^{\ell})+2\epsilon\phi\mathbf{d}(A).$
@@
$\delta_{G[A]}S_{t}$ $\leq\delta_{H}S-\phi\mathbf{d}(A_{t}^{\ell}\setminus S_{t})$
$\leq(1+\epsilon)\phi\mathbf{d}(A_{t}^{\ell})+2\epsilon\phi\mathbf{d}(A)-\phi\mathbf{d}(A_{t}^{\ell}\setminus S_{t})$
$=\phi\mathbf{d}(S_{t}\cap A_{t}^{\ell})+\epsilon\phi\mathbf{d}(A_{t}^{\ell})+2\epsilon\phi\mathbf{d}(A)$
@@
concluding the first statement. For the second statement, observe that $\delta_{H}S\leq(1+\epsilon)\mathbf{s}(A)\leq(1+\epsilon)\cdot 2\phi\mathbf{d}(A)$ and $\delta_{H}S\geq c_{H}(S,\{t\})=12\phi\mathbf{d}(S_{t})$, so
$\mathbf{d}(S_{t})\leq\frac{\delta_{H}S}{12\phi}\leq\frac{(1+\epsilon)\cdot 2\phi\mathbf{d}(A)}{12\phi}\leq\frac{\mathbf{d}(A)}{3},$
concluding the second statement and the proof. ∎
The algorithm defines the flow $f_{t}$ as follows. First, scale the flow $f$ by factor $1/(12\phi)$ and restrict the flow to edges in $G[A\setminus S_{t}]$. In other words, the flow on edges incident to $S_{t}\cup\{s,t,x\}$ are removed. Then, decompose the flow $f_{t}$ into paths (using, for example, a dynamic tree [ST83]). For each vertex $v\in A_{t}^{\ell}\setminus S_{t}$, remove enough paths starting at $v$ until it is the start of exactly $\mathbf{d}(v)/12$ total capacity of paths, and for each vertex $v\notin A_{t}^{\ell}\setminus S_{t}$, remove all paths starting at $v$; let the remaining flow be $f_{t}$. The claim below shows that this last step is always possible, and that $f_{t}$ satisfies the condition below as required by step ii.
@@
$f_{t}$ is a flow from $A_{t}^{\ell}\setminus S_{t}$ to $A_{t}^{r}\setminus S_{t}$ such that each vertex $v\in A_{t}^{\ell}\setminus S_{t}$ is the source of exactly $\mathbf{d}(v)/12$ flow, and each vertex $v\in A_{t}^{r}\setminus S_{t}$ is the sink of at most $\mathbf{d}(v)$ flow. The flow $f$ has congestion $1/(12\phi)$.
@@
    Each vertex $v\in A_{t}^{\ell}\setminus S_{t}$ is the source of at least $\mathbf{d}(v)/12$ flow, and

### v3

@@ 5.2 Fair Cut/Flow Algorithm
Given a congestion-approximator, the most convenient min-cut/max-flow algorithm is the fair cut/flow algorithm of [LNPS23]. We state the definition of a fair cut/flow and then cite the main result of [LNPS23].
@@
Theorem 5.9 (Fair cut/flow algorithm [LNPS23]).
@@ 5.3 Cut-Matching Game and Trimming
We follow the cut-matching game treatment in [SW19]: either find a “balanced” cut of small capacity, or ensure that a “large” part of the graph mixes with low congestion. The following lemma is similar to Theorem 2.2 of [SW19] with one key difference: there is no built-in flow subroutine, so the algorithm makes black-box calls to the fair cut/flow algorithm of Theorem 5.10.
In past work [RST14, AKL+24], the analysis of the cut-matching game for capacitated graphs has only been sketched, referencing the fact that a capacitated graph can be modelled by an uncapacitated graph with parallel edges (at a cost). For completeness, we provide a full proof of this capacitated case in Appendix A.
@@
Property (3) asserts that either an approximately balanced cut is found, or the vertex weighting $\mathbf{d}|_{A\setminus R}$ mixes with low congestion (with high probability). In our algorithm for Theorem 5.1, we actually want the weighting $\mathbf{d}_{A\setminus R}+\deg_{\partial R}$ to mix in the second case. To guarantee this stronger property, we augment the set $R$ into $R\cup B$ through one additional call to the fair cut/flow algorithm of Theorem 5.10. The algorithm is similar to the flow-based expander trimming procedure in [SW19]. For completeness, we defer the algorithm and proof to Appendix B. Note that the setting, including assumption 1, is the same as Theorem 5.11.
@@ 5.4 Clustering Algorithm
For any recursive call $A^{\prime}\subseteq A$, we have $\mathbf{d}^{\prime}(A^{\prime})\leq(1-\frac{1}{36T})\mathbf{d}(A)$ where $\mathbf{d}^{\prime}=\textup{deg}_{\partial\mathcal{P}_{L}\cup\partial A^{\prime}}|_{A^{\prime}}$.
@@
$\leq\frac{7}{6}\cdot\frac{2}{3}\mathbf{d}(A)+\frac{1}{36}\mathbf{d}(A)+\frac{1}{6}\mathbf{d}(A)$
$=\frac{35}{36}\mathbf{d}(A)$
$\leq\left(1-\frac{1}{36T}\right)\mathbf{d}(A),$
@@
For a given recursive call $A$, define its recursion depth inductively as follows: the initial call $A\leftarrow V$ has depth 00, and given a recursive call $A$ of depth $d$, all of its recursive calls have depth $d+1$. By Claim 5.14, the value of $\mathbf{d}(A)$ decreases multiplicatively by factor $1/(36T)$ on each recursive call, so the maximum recursion depth is $O(T\log(nW))$.
@@ 6 Approximate Maximum Flow
From Theorem 5.1 and Claim 5.2, we obtain an algorithm that constructs a congestion-approximator of quality $O(\log^{10}(nW))$ in $\tilde{O}(m)$ time. Recall that Sherman’s framework [She13, She17] translates a congestion-approximator of quality $\alpha$ to a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}\alpha m)$. Thus, for any parameter $\epsilon>0$, we obtain a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}m)$.
Acknowledgement
JL would like to thank Evangelos Kosinas for pointing out an error in the initial version of Lemma 4.7.
@@ Appendix A Cut-Matching Game
Our setup resembles Appendix B of [SW19] with a few minor changes. In particular, since our flow routine is more restrictive, we have to adapt the algorithm to handle our flow outputs.
We begin with notation from [SW19]. For simplicity, we avoid working with the subdivision graph in [RST14, SW19]. Define a $A$-commodity flow as a multi-commodity flow where each vertex $v\in A$ is the source of quantity $\mathbf{d}(v)$ of its distinct flow commodity. Only for analysis, we consider a $A\times A$ flow-matrix $\mathbf{F}\in\mathbb{R}^{A\times A}_{\geq 0}$ which encodes information about a $A$-commodity flow. We say that $\mathbf{F}$ is routable with congestion $c$ if there exists a $A$-commodity flow $f$ such that, simultaneously for all $u,v\in A$, we have that $u$ can send quantity $\mathbf{F}(u,v)$ of its own commodity to $v$, and the amount of flow through each edge is at most $c$.
The algorithm initializes flow-matrix $\mathbf{F}_{0}\in\mathbb{R}^{A\times A}_{\geq 0}$ as the diagonal matrix with value $\mathbf{d}(v)$ on entry $\mathbf{F}(v,v)$. Trivially, $\mathbf{F}$ is routable with zero congestion. The algorithm initializes $A_{0}=A$ and $R_{0}=\emptyset$, and then proceeds for at most $T=O(\log^{2}(nW))$ rounds. For each round $t\in[T]$, the algorithm implicitly updates $\mathbf{F}_{t-1}$ to $\mathbf{F}_{t}$ such that it is routable with congestion $t/\phi$. The operation for implicitly updating $\mathbf{F}_{t-1}$ will be described explicitly later on, but we ensure that row sums do not change from $\mathbf{F}_{t-1}$ to $\mathbf{F}_{t}$, i.e., there is always $\mathbf{d}(v)$ total quantity of each commodity $v\in A$ spread among the vertices. For each round $t\in[T]$, the algorithm (explicitly) finds a partition of $A_{t-1}$ into $A_{t}^{\ell},A_{t}^{r}$, and then computes
@@
    A (possibly empty) set $S_{t}\subseteq A$ satisfying $\delta_{G[A]}S_{t}\leq\phi\mathbf{d}(S\cap A_{t-1})+\frac{\phi}{6T^{2}}\mathbf{d}(A)$ and $\mathbf{d}(S_{t})\leq 0.6\mathbf{d}(A)$, and
@@
    A (possibly empty) flow $f_{t}$ from $A_{t}^{\ell}\setminus S_{t}$ to $A_{t}^{r}\setminus S_{t}$ such that each vertex $v\in A_{t}^{\ell}\setminus S_{t}$ is the source of exactly $\mathbf{d}(v)/24$ flow, and each vertex $v\in A_{t}^{r}\setminus S_{t}$ is the sink of at most $\mathbf{d}(v)$ flow. The flow $f$ has congestion $1/\phi$.
@@
For any round $t$, we have $\delta_{G[A]}R_{t}\leq\phi\mathbf{d}(R_{t})+\frac{\phi}{6T}\mathbf{d}(A)$ and $\mathbf{d}(R_{t})\leq 0.6\mathbf{d}(A)$.
@@
For the second statement, consider the round $t$ with $\mathbf{d}(R_{t})\geq\mathbf{d}(A)/(6T)$, if it exists, at which point the algorithm terminates. (If there is no such $t$, then we are done.) We have $\mathbf{d}(R_{t-1})<\mathbf{d}(A)/(6T)$ and $\mathbf{d}(R_{t-1}\cup S_{t})=\mathbf{d}(R_{t})\geq\mathbf{d}(A)/(6T)$, and the final set $S_{t}$ satisfies $\mathbf{d}(S_{t})\leq 0.6\mathbf{d}(A_{t-1})\leq 0.6\mathbf{d}(A)$. It follows that the new set $R_{t}$ satisfies
$\mathbf{d}(R_{t})=\mathbf{d}(R_{t-1}\cup S_{t})\leq\mathbf{d}(R_{t-1})+\mathbf{d}(S_{t})\leq\frac{\mathbf{d}(A)}{6T}+0.6\mathbf{d}(A)\leq\frac{2}{3}\mathbf{d}(A)$
for large enough $T$, concluding the second statement and the proof. ∎
@@
$\boldsymbol{\mu}_{t}=\frac{\sum_{u\in A_{t}}\mathbf{F}_{t}(u)}{\mathbf{d}(A_{t})}=\arg\min_{\boldsymbol{\mu}(v)\in\mathbb{R}^{A}}\sum_{u\in A_{t}}\mathbf{d}(u)\left\lVert\frac{\mathbf{F}_{t}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}\right\rVert_{2}^{2}$ (2)
is the weighted average of the flow-vectors in $A_{t}$, which is also the minimizer of $\psi(t)$ when treated as a function of $\boldsymbol{\mu}$. The latter fact can be verified separately for each coordinate $v\in A$; namely,
@@ A.2 Computing the Set and Flow
In this subsection, we describe a round of the algorithm in detail, following Lemma B.3 of [SW19] with some minor changes. Recall that on each iteration $t\in[T]$, the algorithm first computes a partition of $A_{t-1}$ into $A_{t}^{\ell},A_{t}^{r}$, and then computes a set $S_{t}$ and flow $f_{t}$ satisfying certain properties. The algorithm then updates $A_{t}\leftarrow A_{t-1}\setminus S_{t}$ and $R_{t}\leftarrow R_{t-1}\cup S_{t}$. The choice of partition $A_{t}^{\ell},A_{t}^{r}$ will depend on an implicitly represented flow-matrix $\mathbf{F}_{t}$ that is useful for the analysis.
@@ A.2.1 Constructing the partition
We first cite a lemma of [RST14] whose proof translates directly to the setting with vertex weighting $\mathbf{d}$.
Lemma A.5 (Lemma 3.3 of [RST14]).
@@
    $\mathbf{d}(A_{t}^{\ell})\leq\mathbf{d}(A_{t-1})/8$ and $\mathbf{d}(A_{t}^{r})\geq\mathbf{d}(A_{t-1})/2$,
@@
    $(\mathbf{p}(v)-\eta)^{2}\geq\frac{1}{9}(\mathbf{p}(v)-\bar{\mu}_{t-1})^{2}$ for each vertex $v\in A^{\ell}_{t}$, and
  4. (d)
    $\sum_{v\in A_{t}^{\ell}}\mathbf{d}(v)\cdot(\mathbf{p}(v)-\bar{\mu}_{t-1})^{2}\geq\frac{1}{80}\sum_{v\in A_{t-1}}\mathbf{d}(v)\cdot(\mathbf{p}(v)-\bar{\mu}_{t-1})^{2}$.
We conclude this section with a lemma about the behavior of projections onto a random unit vector.
Lemma A.6 (Lemma 3.4 of [KRV09a]).
@@
and
$(\mathbf{p}(u)-\bar{\mu}_{t-1})^{2}\leq\frac{c\log n}{n}\left\lVert\frac{\mathbf{F}_{t-1}(u)}{\mathbf{d}(u)}-\boldsymbol{\mu}_{t-1}\right\rVert_{2}^{2},$
each with probability at least $1-n^{-c/4}$.
@@ A.2.2 Max-flow/min-cut call
$A\leftarrow A,\,\epsilon\leftarrow\frac{1}{18T^{2}},\,\gamma\leftarrow\frac{\epsilon\phi}{2},\,\beta\leftarrow\max\{1,(12\phi+\epsilon\gamma)(\kappa+2)\},\,\,\mathbf{s}\leftarrow\phi\mathbf{d}|_{A_{t-1}\setminus A_{t}^{r}}+\epsilon\phi\mathbf{d},\,\mathbf{t}\leftarrow 12\phi\mathbf{d}|_{A_{t}^{r}},$ (3)
@@
From Theorem 5.10, we obtain an $(1+\epsilon)$-approximate fair cut/flow pair $(S,f)$. The algorithm sets $S_{t}\leftarrow S\setminus\{s,x\}$, which satisfies the condition below as required by step B.3.
@@
$\delta_{G[A]}S_{t}\leq\phi\mathbf{d}(S\cap A_{t-1})+3\epsilon\phi\mathbf{d}(A)$ and $\mathbf{d}(S_{t})\leq 0.6\mathbf{d}(A)$,
@@
    The edges incident to $s$, which have total capacity
    $c_{H}(\{s\},V_{H}\setminus S)=\mathbf{s}(A\setminus S_{t})\geq\phi\mathbf{d}((A_{t-1}\setminus A_{t}^{r})\setminus S_{t}),$
@@
$\delta_{H}S\geq\delta_{G[A]}S_{t}+\phi\mathbf{d}((A_{t-1}\setminus A_{t}^{r})\setminus S_{t}).$
@@
$\delta_{H}S\leq(1+\epsilon)\mathbf{s}(A)=(1+\epsilon)(\phi\mathbf{d}(A_{t-1}\setminus A_{t}^{r})+\epsilon\phi\mathbf{d}(A))\leq(1+\epsilon)\phi\mathbf{d}(A_{t-1}\setminus A_{t}^{r})+2\epsilon\phi\mathbf{d}(A).$
@@
$\delta_{G[A]}S_{t}$ $\leq\delta_{H}S-\phi\mathbf{d}((A_{t-1}\setminus A_{t}^{r})\setminus S_{t})$
$\leq(1+\epsilon)\phi\mathbf{d}(A_{t-1}\setminus A_{t}^{r})+2\epsilon\phi\mathbf{d}(A)-\phi\mathbf{d}((A_{t-1}\setminus A_{t}^{r})\setminus S_{t})$
$=\phi\mathbf{d}(S_{t}\cap(A_{t-1}\setminus A_{t}^{r}))+\epsilon\phi\mathbf{d}(A_{t-1}\setminus A_{t}^{r})+2\epsilon\phi\mathbf{d}(A)$
@@
concluding the first statement. For the second statement, observe that $\delta_{H}S\leq(1+\epsilon)\mathbf{s}(A)\leq(1+\epsilon)^{2}\phi\mathbf{d}(A)$ and $\delta_{H}S\geq c_{H}(S,\{t\})=12\phi\mathbf{d}(S_{t}\cap A_{t}^{r})$, so
$\mathbf{d}(S_{t}\cap A_{t}^{r})\leq\frac{\delta_{H}S}{12\phi}\leq\frac{(1+\epsilon)^{2}\phi\mathbf{d}(A)}{12\phi}\leq\frac{\mathbf{d}(A)}{10}$
for $\epsilon>0$ small enough. By property b of Lemma A.5, we have $\mathbf{d}(A_{t}^{r})\geq\mathbf{d}(A_{t-1})/2$, so $\mathbf{d}(S_{t})\leq\mathbf{d}(S_{t}\cap A_{t}^{r})+\mathbf{d}(A_{t-1}\setminus A_{t}^{r})\leq\mathbf{d}(A)/10+\mathbf{d}(A_{t-1})/2\leq 0.6\mathbf{d}(A)$, concluding the second statement and the proof. ∎
The algorithm defines the flow $f_{t}$ as follows. First, scale the flow $f$ by factor $1/(12\phi)$ and restrict the flow to edges in $G[A\setminus S_{t}]$. In other words, the flow on edges incident to $S_{t}\cup\{s,t,x\}$ are removed. Then, decompose the flow $f_{t}$ into paths (using, for example, a dynamic tree [ST81]). For each vertex $v\in A_{t}^{\ell}\setminus S_{t}$, remove enough paths starting at $v$ until it is the start of exactly $\mathbf{d}(v)/24$ total capacity of paths, and for each vertex $v\notin A_{t}^{\ell}\setminus S_{t}$, remove all paths starting at $v$; let the remaining flow be $f_{t}$. The claim below shows that this flow removal step is always possible, and that $f_{t}$ satisfies the condition below as required by step ii.
@@
$f_{t}$ is a flow from $A_{t}^{\ell}\setminus S_{t}$ to $A_{t}^{r}\setminus S_{t}$ such that each vertex $v\in A_{t}^{\ell}\setminus S_{t}$ is the source of exactly $\mathbf{d}(v)/24$ flow, and each vertex $v\in A_{t}^{r}\setminus S_{t}$ is the sink of at most $\mathbf{d}(v)$ flow. The flow $f$ has congestion $1/(12\phi)$.
@@
    Each vertex $v\in A_{t}^{\ell}\setminus S_{t}$ is the source of at least $\mathbf{d}(v)/24$ flow, and